# Estimating Body and Hand Motion in an Ego-sensed World

### v2

@@ Abstract
We present EgoAllo, a system for human motion estimation from a head-mounted device. Using only egocentric SLAM poses and images, EgoAllo guides sampling from a conditional diffusion model to estimate 3D body pose, height, and hand parameters that capture a device wearer’s actions in the allocentric coordinate frame of the scene. To achieve this, our key insight is in representation: we propose spatial and temporal invariance criteria for improving model performance, from which we derive a head motion conditioning parameterization that improves estimation by up to 18%. We also show how the bodies estimated by our system can improve the hands: the resulting kinematic and temporal constraints can reduce errors in noisy monocular estimates by 40%. Project page: https://egoallo.github.io
@@ 1 Introduction
Head-mounted devices are permeating the mainstream market, with applications in areas like virtual reality, content creation, and assistive technologies. For research in machine perception, these devices also present exciting opportunities in the form of egocentric sensing. Sensors from wearable devices provide abundant observations of 3D environments, while capturing the embodied perspective of human agents as they navigate and interact with the world around them.
What can we understand from these devices as their adoption widens? As a starting point, parallax-inducing egomotion provides excellent conditions for advances in 3D reconstruction and scene understanding [109, 17, 65]. Limiting perception to only the surrounding world, however, would neglect a crucial piece of the ego-sensory puzzle: the individual whose decisions shape the inputs. Capturing the wearer’s actions and motion in addition to the scene promises to unlock applications across augmented and virtual reality, robotics, and general human behavior analysis, while unveiling action-tied semantics in the scene itself.
We therefore introduce EgoAllo, a system that uses egocentric inputs to estimate the wearer’s actions in the allocentric frame of the world. This is a difficult task: while body parts like hands occasionally appear in egocentric image frames, most body parameters are never directly observed by these devices. Accurate estimates in the global frame also require harmony between both pose and height parameters. These must be consistent with the egomotion and scene scale, aligning estimated feet to the ground and head to the sensed camera height.
We cast estimation as guided sampling from a human motion diffusion model. We consider two inputs: head poses from SLAM and visual hand observations from images. We then estimate body pose, height, and hand parameters, unlike most prior works in egocentric human motion estimation [46, 6] that focus on body pose.
We achieve this by training a head motion-conditioned diffusion model as a motion prior, and guiding its sampling to recover a hand-body sequence that aligns with image observations. Our results are enabled by a key insight: that the representation used for head pose conditioning is critical for accurate ego-sensed motion estimation. We study choices for this representation by (1) identifying desirable spatial and temporal invariance properties that are not fulfilled by existing systems, (2) using these properties to derive improved parameterizations for our motion prior, and (3) presenting a systematic evaluation that shows between 4.9% and 17.9% error difference when compared against prior work. Furthermore, we show how the resulting system can improve hand estimation, reducing errors by over 40% compared to noisy monocular estimates.
@@ 2 Related Work
Priors for human motion. The primary challenge of ego-sensed human motion estimation is limited observability; a prior is required to resolve ambiguities. For human motion, these priors are typically framed as unconditional distributions over plausible human motions. These distributions can be represented either by modeling the physical constraints of our world [70, 48, 64, 5] or by learning generative models of human motion directly from data. For learning unconditional priors, classical data-driven approaches include fitting mixtures-of-Gaussians to 3D keypoint trajectories [24], while modern approaches include training variational autoencoders [36, 72] to model either autoregressive transitions [71, 49, 14] or full spatiotemporal sequences [21]. After training, these priors can be applied to estimation problems in iterative optimization frameworks [99, 39, 71]. EgoAllo is built on the same intuition as these methods, but follows previous work in ego-sensed motion estimation and uses a task-specific conditional prior.
Figure 2: Overview of components in EgoAllo. We restrict the diffusion model to local body parameters $\{\Theta^{t},\beta^{t}\}_{t},\{\psi_{j}^{t}\}_{t,j}$ (Section 3.1.1). An invariant parameterization (Section 3.1.2) of SLAM poses is used to condition a diffusion model. These can be placed into the global coordinate frame via global alignment (Section 3.2.1) to input poses. When available, egocentric video is used for hand detection via HaMeR [63], which can be incorporated into samples via guidance (Section 3.2.2).
@@
Human motion from egocentric observations. EgoAllo builds on intuition from several prior works in egocentric sensing for human motion estimation. Many rely on fisheye cameras that place the wearer’s body into the field of view [26, 89, 88, 92, 95, 73, 88, 93]. Other approaches rely on body-mounted cameras [80], simulation-based physical plausibility [104, 105, 52], body- and hand-mounted inertial sensors [102, 103, 44], handheld controllers [6, 27, 28], and interaction cues from other humans [56]. Concurrent works have also used the Nymeria [53] dataset for egocentric motion with language description outputs [23], as well for online settings with scene geometry and CLIP [67] feature inputs [20]. Most relevantly, EgoEgo [46] demonstrates how human body poses can be estimated offline without body observability assumptions. The authors accomplish this by carefully integrating several components: a monocular SLAM system [86], a pose-conditioned gravity vector regression network, an optical flow feature-conditioned head orientation and scale regression network, and a head pose-conditioned body diffusion model. EgoAllo differs in both inputs—we study conditioning parameters computed from the metric SLAM poses provided by devices like Project Aria [82]—and outputs—we consider body height variation and hand poses.
Conditioning for ego-sensed poses. Prior works vary in how head pose information is parameterized and used as neural network input. AvatarPoser [27] and BoDiffusion [6] parameterize head pose as four components: world-frame orientation, orientation deltas, world-frame position, and world-frame position deltas. These works are focused on settings with VR controller input, and parameterize controller pose inputs the same way. EgoEgo [46]’s diffusion model uses only absolute head positions and orientations, but, similar to HuMoR [71], in implementation defines a per-sequence canonical coordinate frame to ensure that all input trajectories passed to the model are aligned with the same initial $xy$ position and forward direction. In our work, we will refer to this as sequence canonicalization. Finally, EgoPoser [28] proposes a similar scheme that aligns initial positions for both head pose and controller pose inputs. We propose an alternative to these parameterizations that is motivated by the robustness and generalization benefits of invariance, as observed in prior work for designing both representations [94, 51, 78, 9, 113, 101, 108] and neural network architectures [43, 31, 107, 87, 8, 7, 11, 13, 97, 98]. Specifically, we introduce in Section 3.1.2 a parameterization that is invariant to both spatial and temporal shifts.
@@ 3 Method
We study the problem of using sensors from an egocentric device to estimate the actions of the wearer in an allocentric coordinate frame. We assume a flat floor and two egocentric inputs: poses from the device’s SLAM system and camera images.
@@ 3.1.1 Diffusion output representation
As output, we sample joint rotations, body shapes, and binary contact predictions $\vec{x}_{0}^{t}=\{\Theta^{t},\beta^{t},\psi_{j=1\dots 21}^{t}\}$, where body shape $\beta^{t}$ is supervised to be equal for all timesteps and $\psi_{j}^{t}$ is a per-joint contact indicator. Notably, these parameters are all local—we discuss how outputs can be placed into the allocentric coordinate frame in Section 3.2.1.
@@ 3.1.2 Invariant conditioning
To encode absolute height, we assume that the world frame’s +$z$-axis faces upwards, and that the ground is located at $z=0$. Ground parameters are directly available in the training data [54]; at test time, we can also extract these parameters from sparse SLAM points via RANSAC (Appendix A.1).
@@
Relative transforms alone, however, do not encode information relative to the scene or floor: full trajectories can even be flipped upside down without impacting $\Delta\mathbf{T}_{\text{cpf}}^{t-1,t}$. We therefore propose to ground relative motion to the floor plane with a transformation between the CPF frame and a per-timestep canonical frame, which is computed by projecting the CPF frame to the floor. This encodes head height and orientation. Our full representation then becomes:
@@
We visualize an example of a canonical frame in Figure 3. Canonical frames are positioned by projecting the CPF origin to the floor plane; given standard bases $\mathbf{e}_{\{x,y,z\}}$, we compute:
@@
$\mathbf{R}_{\text{world},\text{canonical}}^{t}$ $=\mathbf{R}_{z}\left(-\text{arctan2}\left(\mathbf{e}_{x}^{\top}\vec{v}^{\ t},\mathbf{e}_{y}^{\top}\vec{v}^{\ t}\right)\right)\mathbf{R}_{\text{world},\text{cpf}}^{t}.$ (7)
@@ 3.2 Estimation via sampling
We use our local body representation and invariant conditioning strategies to train a motion prior in the form a denoising diffusion model [22]. Given diffusion step $n=N\dots 1$, we follow [69] and approximate the denoising process as:
@@ 3.2.1 Global alignment
where $\mathbf{T}_{\text{cpf},\text{root}}^{(\Theta^{t},\beta^{t})}$ computes the transform between the root of the human and their CPF frame for a given set of local pose and shape parameters. Similar processes are applied in [28, 27, 6]. In contrast to directly outputting absolute body transformations from the diffusion model [46], this guarantees exact alignment between estimates and the input SLAM sequences.
@@ 3.2.2 Guidance losses
Our diffusion model learns a distribution of human motion conditioned on the central pupil frame motion. At test time, we incorporate constraints from physical priors and visual hand observations via guidance [112, 29, 10]. Similar to [45, 34], we accomplish this by applying losses to the joint rotations $\Theta=\{\Theta^{1},\dots,\Theta^{T}\}$ predicted by $\mu_{\theta}(\vec{x}_{n},n,\vec{c})$. We treat the body shape $\beta^{t}$ and contacts $\psi^{t}_{j=1\dots 21}$ as fixed and optimize over body and finger pose to minimize hand observation, skating, and prior costs with a Levenberg-Marquardt optimizer:
@@
The 3D objective $\mathcal{E}_{\text{hands3D}}^{(\Theta)}$ minimizes the distance between the detected hand parameters and the corresponding SMPL-H hand parameters, in terms of wrist pose and local joint rotations. With $\Pi_{K}$ as projection with camera intrinsics $K$, $\mathbf{p}_{world,j}^{(\Theta^{t})}\in\mathbb{R}^{3}$ as the world position for joint $j$ at time $t$, and $\mathbf{T}_{\text{camera},\text{cpf}}$ from the device calibration, the reprojection loss is:
@@
To reduce foot skating, we use contact predictions to apply a skating loss [71, 99] for each time $t$ and joint $j$:
@@
Finally, we minimize a prior term $\mathcal{E}^{(\Theta)}_{\text{prior}}$. This term penalizes deviations between each constrained rotation matrix $\Theta^{t}$ and the original output rotations $\hat{\Theta}^{t}$ from our denoiser $\mu_{\theta}(\vec{x}_{n},n,\vec{c})$, in terms of both joint rotation and rotational velocity.
@@ 4.1 Body estimation
In our first set of experiments, we evaluate body estimation from only device SLAM poses, without considering images or hands. This setting allows us to isolate the advantages of our body motion prior, while facilitating direct comparison against methods that do not consider hands.

### v3

@@ Abstract
We present EgoAllo, a system for human motion estimation from a head-mounted device. Using only egocentric SLAM poses and images, EgoAllo guides sampling from a conditional diffusion model to estimate 3D body pose, height, and hand parameters that capture a device wearer’s actions in the allocentric coordinate frame of the scene. To achieve this, our key insight is in representation: we propose spatial and temporal invariance criteria for improving model performance, from which we derive a head motion conditioning parameterization that improves estimation by up to 18%. We also show how the bodies estimated by our system can improve hand estimation: the resulting kinematic and temporal constraints can reduce world-frame errors in single-frame estimates by 40%.
@@ 1 Introduction
Head-mounted devices are becoming increasingly mainstream. In addition to offering new challenges for 3D scene understanding [111, 18, 68, 57], egocentric sensors from these devices are unique in that their outputs are coupled to a human wearer’s motion in the world. Using these sensors to understand the wearer in addition to the scene around them is essential for applications in augmented reality, robotics, and assistive technologies.
We therefore introduce EgoAllo, a system that uses egocentric inputs to estimate the wearer and their motion in the world, or allocentric, coordinate frame. We take as input sensed metric SLAM head poses and egocentric video from devices like Project Aria [61]. We then estimate as output human body pose, height, and hand motion parameters.
This is a difficult task: while body parts like hands occasionally appear in egocentric frames, most body parameters are never directly observed. To ensure that estimates are consistent with both the scene and sensed egomotion, harmony is also required between pose and height parameters. This setting differs from most prior works in egocentric human motion estimation [48, 7], which focus on body pose and do not address the challenges of height and hand motion.
Our proposed system uses a head pose-conditioned diffusion model as a motion prior, as well as a Levenberg-Marquardt guidance optimizer for sampling hand-body sequence that align with image observations. Our results are enabled by a key insight: that the representation used for head pose conditioning is critical for accurate full-body motion estimation. We study choices for this representation by (1) identifying desirable spatial and temporal invariance properties that are not fulfilled by existing systems and (2) using these properties to derive improved parameterizations for our motion prior.
We systematically evaluate our system on four datasets. For body estimation, we find that improving the conditioning parameterization leads to an accuracy improvement between 4.9% and 17.9%. Furthermore, we observe that the resulting system can improve hand estimation, reducing world-frame errors by over 40% compared to single-frame estimates. Code, model, and more results can be found on our project webpage.
@@ 2 Related Work
Priors for human motion. The primary challenge of ego-sensed human motion estimation is limited observability; a prior is required to resolve ambiguities. For human motion, these priors are typically framed as unconditional distributions over plausible human motion. Distributions can be represented either by modeling the physical constraints of our world [73, 50, 67, 6] or by learning generative models of human motion directly from data. For learning unconditional priors, classical data-driven approaches include fitting mixtures-of-Gaussians to 3D keypoint trajectories [25], while modern approaches include training variational autoencoders [38, 75] to model either autoregressive transitions [74, 51, 15] or full spatiotemporal sequences [22]. After training, these priors can be applied to estimation problems in iterative optimization frameworks [102, 41, 74]. EgoAllo is built on the same intuition as these methods, but follows previous work in ego-sensed motion estimation and uses a task-specific conditional prior.
Figure 2: Overview of components in EgoAllo. We restrict the diffusion model to local body parameters (Section 3.1.1). An invariant parameterization $g(\cdot)$ (Section 3.1.2) of SLAM poses is used to condition a diffusion model. These can be placed into the global coordinate frame via global alignment (Section 3.2.1) to input poses. When available, egocentric video is used for hand detection via HaMeR [66], which can be incorporated into samples via guidance (Section 3.2.2).
@@
Human motion from egocentric observations. EgoAllo builds on intuition from several prior works in egocentric sensing for human motion estimation. Many rely on fisheye cameras that place the wearer’s body into the field of view [27, 92, 91, 95, 98, 76, 91, 96]. Other approaches rely on body-mounted cameras [83], simulation-based physical plausibility [106, 107, 54], body- and hand-mounted inertial sensors [104, 105, 46, 37], handheld controllers [7, 28, 29], and interaction cues from other humans [59]. Concurrent works have also used the Nymeria [55] dataset for egocentric motion with language description outputs [24], as well for online settings with scene geometry and CLIP [70] feature inputs [21]. Most relevantly, EgoEgo [48] demonstrates how human body poses can be estimated offline without body observability assumptions. The authors accomplish this by carefully integrating several components: a monocular SLAM system [89], a pose-conditioned gravity vector regression network, an optical flow feature-conditioned head orientation and scale regression network, and a head pose-conditioned body diffusion model. EgoAllo differs in both inputs—we study conditioning parameters computed from the metric SLAM poses provided by devices like Project Aria [85]—and outputs—we consider body height variation and hand poses.
Conditioning for ego-sensed poses. Prior works vary in how head pose information is parameterized and used as neural network input. AvatarPoser [28] and BoDiffusion [7] parameterize head pose as four components: world-frame orientation, orientation deltas, world-frame position, and world-frame position deltas. These works focus on settings with VR controller input, and parameterize controller pose inputs the same way. EgoEgo [48]’s diffusion model uses only absolute head positions and orientations, but similar to HuMoR [74], in implementation defines a per-sequence canonical coordinate frame to ensure that all input trajectories passed to the model are aligned with the same initial $xy$ position and forward direction. In our work, we refer to this as sequence canonicalization. Finally, EgoPoser [29] proposes a similar scheme that aligns initial positions for both head pose and controller pose inputs. We propose an alternative to these parameterizations that is motivated by the robustness and generalization benefits of invariance, as observed in prior work for designing both representations [97, 53, 81, 10, 115, 103, 110] and neural network architectures [45, 32, 109, 90, 9, 8, 12, 14, 100, 101]. Specifically, we introduce in Section 3.1.2 a parameterization that is invariant to both spatial and temporal shifts.
@@ 3 Method
We study the problem of using sensors from an egocentric device to estimate the actions of the wearer in an allocentric coordinate frame. We assume a flat floor and two inputs: poses from the device’s SLAM system and egocentric video.
@@ 3.1.1 Diffusion output representation
As output, we sample body and hand joint rotations, body shapes, and binary contact predictions $\vec{x}_{0}^{t}=\{\Theta^{t},\beta^{t},\psi_{j=1\dots 21}^{t}\}$, where body shape $\beta^{t}$ is supervised to be equal for all timesteps and $\psi_{j}^{t}$ is a per-joint contact indicator. Notably, these parameters are all local—we discuss how outputs can be placed into the allocentric coordinate frame in Section 3.2.1.
@@ 3.1.2 Invariant conditioning
To encode absolute height, we assume that the world frame’s +$z$-axis faces upwards, and that the ground is located at $z=0$. Ground parameters are directly available in the training data [56]; at test time, we can also extract these parameters from sparse SLAM points via RANSAC (Appendix A.3.3).
@@
Relative transforms alone do not encode information relative to the scene or floor: full trajectories can even be flipped upside down without impacting $\Delta\mathbf{T}_{\text{cpf}}^{t-1,t}$. We therefore propose to ground relative motion to the floor plane with a transformation between the CPF frame and a per-timestep canonical frame, which is computed by projecting the CPF frame to the floor. This encodes head height and orientation. Our full representation then becomes:
@@
We visualize an example of a canonical frame in Figure 3 and our full representation in Appendix A.1. Canonical frames are positioned by projecting the CPF origin to the floor plane; given standard bases $\mathbf{e}_{\{x,y,z\}}$, we compute:
@@
$\mathbf{R}_{\text{world},\text{canonical}}^{t}$ $=\mathbf{R}_{z}\left(-\text{arctan2}\left(\mathbf{e}_{x}^{\top}\vec{v}^{\ t},\mathbf{e}_{y}^{\top}\vec{v}^{\ t}\right)\right).$ (7)
@@ 3.2 Estimation via sampling
We use our local body representation and invariant conditioning strategies to train a motion prior in the form of a denoising diffusion model [23]. Given diffusion step $n=N\dots 1$, we follow [72] and approximate the denoising process as:
@@ 3.2.1 Global alignment
where $\mathbf{T}_{\text{cpf},\text{root}}^{(\Theta^{t},\beta^{t})}$ computes the transform between the root of the human and their CPF frame for a given set of local pose and shape parameters. Similar processes are applied in [29, 28, 7]. In contrast to directly outputting absolute body transformations from the diffusion model [48], this guarantees exact alignment between estimates and input SLAM sequences.
@@ 3.2.2 Guidance losses
Our diffusion model learns a distribution of human motion conditioned on the central pupil frame motion. At test time, we incorporate constraints from physical priors and visual hand observations via guidance [114, 30, 11]. Similar to [47, 35], we accomplish this by applying costs to the joint rotations $\Theta=\{\Theta^{1},\dots,\Theta^{T}\}$ predicted by $\mu_{\theta}(\vec{x}_{n},n,\vec{c})$. We treat the body shape $\beta^{t}$ and contacts $\psi^{t}_{j=1\dots 21}$ as fixed and optimize over body and finger pose to minimize hand observation, skating, and prior costs with a Levenberg-Marquardt optimizer:
@@
The 3D objective $\mathcal{E}_{\text{hands3D}}^{(\Theta)}$ minimizes the distance between the detected hand parameters and the corresponding SMPL-H hand parameters, in terms of wrist pose and local joint rotations. With $\Pi_{K}$ as projection with camera intrinsics $K$, $\mathbf{p}_{world,j}^{(\Theta^{t})}\in\mathbb{R}^{3}$ as the world position for joint $j$ at time $t$, and $\mathbf{T}_{\text{camera},\text{cpf}}$ from the device calibration, the reprojection cost is:
@@
To reduce foot skating, we use contact predictions to apply a skating cost [74, 102] for each time $t$ and joint $j$:
@@
Finally, we minimize a prior cost $\mathcal{E}^{(\Theta)}_{\text{prior}}$. This cost penalizes deviations between joint rotations $\Theta^{t}$ and rotations $\hat{\Theta}^{t}$ from the denoiser $\mu_{\theta}(\vec{x}_{n},n,\vec{c})$. We include terms for absolute rotation, rotational velocity, and forward kinematics position.
@@ 4.1 Body estimation
In our first set of experiments, we evaluate body estimation from only device SLAM poses, without considering images or hands. This setting allows us to isolate the advantages of our body motion prior, while directly comparing against methods that do not consider hands.
@@ 4.1.1 Invariant conditioning evaluation
4.1.2 Comparisons against baselines
To further study EgoAllo’s body estimation quality, we compare against three baselines. (1) NoShape. First, NoShape refers to a variation of EgoAllo that turns off shape estimation, and thus cannot estimate the wearer’s height. (2) EgoEgo. We also compare against the human motion diffusion model from EgoEgo [48]. This is similar to EgoAllo, but considers only the SMPL “mean” body shape and uses sequence canonicalized coordinates for conditioning and as model output. (3) VAE+Opt. Finally, we compare against an approach based on the SLAHMR [102] framework for human motion estimation from exocentric video. A key advantage of SLAHMR is that it uses an unconditional motion prior [74] in an optimization framework. It can therefore be adapted to new settings without re-training—we keep the same body pose and shape variables as the original pipeline, but replace the exocentric keypoint [71] cost with an egocentric CPF pose alignment cost.
EgoAllo improves body motion estimates. We report metrics in Table 2 and visualize example outputs in Figure 4. We find that EgoAllo enables significant estimation improvements across all datasets, including accuracy improvements of 20$\sim$30over EgoEgo for both shorter and longer evaluation sequences. We found shape estimation critical for producing metric-scale, grounded estimates of human body motion, with the head aligned to input SLAM poses and the feet planted on the observed ground plane. This is evident in qualitative results, improved grounding metrics, and in the 6$\sim$7MPJPE gap between EgoAllo and the NoShape ablation.
VAE optimization converges poorly. Optimization-based estimation approaches have been effective for settings with keypoint costs [74, 102], but we found convergence difficult in our less constrained setting. In Table 2, we observe poor generalization: VAE+Opt performs competitively on the AMASS test set, but performance deteriorates dramatically when evaluating on RICH or ADT. VAE+Opt outputs in Figure 4 also look overly smoothed, without the same expressiveness as the conditional predictions of EgoAllo or EgoEgo [48]. This highlights the advantage of using a conditional diffusion model problem for this estimation problem.
4.2 Hand estimation
To evaluate hands estimated by EgoAllo, we run HaMeR on the segment of the EgoExo4D [17] validation set that is labeled with 3D hand pose keypoints. We quantitatively compare four hand estimation methods in Table 3. In (1) HaMeR [66], we use HaMeR out-of-the-box on undistorted egocentric RGB images. We do not assume bounding boxes as input; instead, we follow the HaMeR demo code and compute crops using ViTPose [99]. (2) EgoAllo-NoReproj uses all loss terms except for the reprojection loss (Equation 14). Hand guidance is done directly using the 3D wrist poses predicted by HaMeR. (3) EgoAllo-Mono is the same as EgoAllo-NoReproj, but guides hands using the reprojection loss. This accounts for the scale ambiguities that are inherent to the single-frame HaMeR estimates. Finally, (4) EgoAllo-Wrist3D uses both the HaMeR losses and 3D wrist pose losses from Project Aria’s Machine Perception Services [85]—unlike HaMeR, which assumes monocular input, this uses a pair of SLAM cameras that are unique to Project Aria. For fairness across settings, we compute metrics only on timesteps where HaMeR estimates are available.
Results are provided in Table 3. While HaMeR’s local poses (PA-MPJPE) are slightly better, EgoAllo’s hand-body estimation significantly improves how well hands are estimated in the world coordinate system. Compared to HaMeR, EgoAllo-Mono drops MPJPE from $237.90\text{mm}\to 131.45\text{mm}$. Incorporating more accurate wrist pose estimates (EgoAllo-Wrist3D) offers a practical solution for further improvements: $131.45\text{mm}\to 60.08\text{mm}$. Reprojection-based guidance is also important: despite using the same inputs, EgoAllo-NoReproj outputs are worse than EgoAllo-Mono in both MPJPE and PA-MPJPE.
Qualitatively, we observed that high hand estimation errors in naive monocular estimation with HaMeR are explained by a combination of detection failures and monocular ambiguities. Even when detections succeed, the scale and distance of monocular HaMeR estimates are often incorrect or flicker in between frames. Incorporating these hands via guidance with our diffusion motion prior encourages final outputs that obey the kinematic and smoothness constraints imposed by plausible body motion—we provide examples of HaMeR estimates rendered jointly with EgoAllo outputs in Figure 5.
Method MPJPE $\downarrow$ PA-MPJPE $\downarrow$
HaMeR 237.90$\pm$1.89 13.04$\pm$1.89
EgoAllo-NoReproj 143.20$\pm$0.42 14.75$\pm$0.42
EgoAllo-Mono 131.45$\pm$0.39 14.71$\pm$0.39
EgoAllo-Wrist3D 60.08$\pm$0.26 14.38$\pm$0.26
5 Dicussion
Limitations and future work. While the core contributions of EgoAllo are general, the current implementation of our system has a few limitations that we hope to explore in future work. First, diffusion model guidance is a test-time optimization process that depends on hyperparameters and incurs a runtime cost. In the future, it may be possible to bootstrap using outputs from our model to train a feedforward model that avoids this step. Success for hand guidance also still depends on reasonable monocular hand estimates. Estimation can therefore fail as a result of errors like left/right flipping or spurious detections. Finally, we assume flat floors. This is in part because our training data [56] includes floor planes but no detailed scene geometry. As a result, our method will fail in settings like hills or staircases. In the future, we hope to extend our insights to data with more detailed scene information, which concurrent work has highlighted the usefulness of in informing human body estimation [21].
Conclusion. We presented EgoAllo, a system for estimating human motion using sensors from head-mounted devices. EgoAllo jointly estimates human body pose, height, and hand parameters from only egocentric SLAM poses and images. Results highlight the importance of spatial and temporal invariance in conditioning for this problem, while demonstrating how estimated bodies can be used to improve hand estimation.
Supplementary Material
A.1 Invariant Conditioning Visualization
As we observe in Table 1, naively training a model using absolute head poses results in poor estimation performance. The absence of spatial invariance (Invariance 1) explains this result. To visualize this, we show in Figure A.1 two renders of the same human motion trajectory. The second render has the same local body motion as the first, but with the world frame re-defined:
(a) Before
(b) After
Figure A.1: Absolute head pose visualization for a single human motion trajectory, before and after re-defining the world frame.
Because the world frame location is arbitrarily defined, naively conditioning on these poses hinders generalization. Works like EgoPoser [29] have made similar observations.
To fix this, prior works have preprocessed sequences by aligning them to a canonical coordinate frame located at the first timestep of each sequence [48, 74, 21]. However, we observe that this is flawed from the perspective of temporal invariance (invariance 2). To visualize this, we render in Figure A.2 two temporal slices of the same body motion, with one slice starting from the beginning of the motion and another starting from the middle:
(a) First slice
(b) Second slice
Figure A.2: Two slices of the same human motion trajectory.
Next, we consider how the head pose trajectories for each of these slices would look if they were canonicalized by aligning the first timestep. We visualize the resulting head pose trajectories in Figure A.3. Circled in red are four timesteps that are shared between the two slices. Notice that head poses from canonicalized sequences can still differ significantly, even for the same body motion.
(a) First slice
(b) Second slice
Figure A.3: Poses canonicalized by aligning the first timestep.
To achieve both Invariances 1 and 2, EgoAllo’s invariant conditioning paramterization proposes an alternative way to canonicalize head poses. Instead of defining a single canonical coordinate frame for each temporal window, we define a canonical coordinate frame at every timestep. The resulting representation couples relative CPF motion $\Delta\mathbf{T}_{\text{cpf}}^{t}$ with per-timestep canonicalized pose $\mathbf{T}_{\text{canonical, cpf}}^{t}$. These transformations are visualized in Figure A.4. Notice that the transformations that make up this conditioning approach are invariant both to the world coordinate system and to choices in temporal windowing. This enables significant improvements in estimation accuracy (Table 1).
Figure A.4: Transformations that make up the invariant conditioning used by EgoAllo.
A.2 Ancillary Results
A.2.1 Sequence length evaluation
At test time, EgoAllo follows MultiDiffusion [3] for extrapolating to arbitary sequence lengths. To validate this choice, we filter out test sequences shorter than 256 frames and then evaluate both EgoAllo and EgoEgo [48] with subsequences of length 32, 128, and 256. We report MPJPE metrics on these sequences in Table A.1. Both EgoAllo and EgoEgo include windowing strategies for handling longer sequences; unlike prior work, however, we find that accuracy improves even after test set sequence lengths surpass the training set sequence length.
Seqlen 32 128 256
EgoAllo 149.3 130.3 127.9
EgoEgo 187.7 173.8 184.3
Table A.1: Effect of sequence length on MPJPE (mm). EgoAllo is trained with sequences of length 128. EgoEgo [48] is trained with sequences of length 140.
A.2.2 Additional qualitative results
We provide additional qualitative results for the body motion prior in Figure A.5. EgoAllo estimates have the head aligned exactly to input observations and the feet planted realistically on the floor.